OpenStudy (anonymous):

Differential equation: Solve the differential equation by regarding y as the independent variable rather than x. (1+2xy)(dy/dx)= 1+y^2

OpenStudy (anonymous):

So i tried to put it into the format of \[\frac{ dy }{ dx }+P(x)*y=Q(x)\]

OpenStudy (anonymous):

(switching basically the x and y components since y is independent) but i keep getting stuck

OpenStudy (anonymous):

\[1+2xy=(1+y^2)*\frac{ dx }{dy }\] \[\frac{ dx }{ dy }-\frac{ 2xy }{ 1+y^2}=\frac{ 1 }{ 1+y^2}\]

OpenStudy (anonymous):

so then p(y) = \[\frac{ -2y }{ 1+y^2 }\] if i integrate that i get \[-\ln (y^2+1)\] Then to the e i get \[-(y^2+1)\] so then my equation becomes \[-(y^2+1)x'-2y*x=-1\]

OpenStudy (anonymous):

but...i think i did something wrong

OpenStudy (anonymous):

the answer should be(solving through the last equation i wrote) \[=1+y^2)x(y)= \frac{ 1 }{ 2 }[y+(1+y^2)(\tan^{-1} (y)+c)]\]

OpenStudy (anonymous):

ignore the stuff infront of x(y)...typo

OpenStudy (anonymous):

integrating through i get: \[-x(y^2+1)=-y+c\] \[x=\frac{ y+c }{ y^2+1 }\] ..which clearly is wrong

OpenStudy (anonymous):

The integrating factor is almost right. You're correct with \[\mu(y)=e^{-\ln(y^2+1)}\] However, this is not equal to \(-(y^2+1)\): \[e^{-\ln(y^2+1)}=e^{\ln(y^2+1)^{-1}}=\frac{1}{y^2+1}\]

OpenStudy (anonymous):

dang it i did that on an earlier problem too...alright ill try that real quick

OpenStudy (anonymous):

got it, thank you

OpenStudy (anonymous):

Yep, you're welcome